OpenStudy (anonymous):

Find the area of a regular pentagon with side length = 12 and an apothem = 3. Answer 45 90 108 360

OpenStudy (anonymous):

A=(1/2)ap p=60 a=3 A=(1/2)(3)(60) A=90

OpenStudy (anonymous):

can you help me on my next one?